OpenStudy (msmr):

If f(x) has a zero at x=2, 1 and 3, then a factor of f(x) = 3• f(x-2) is...

Parth (parthkohli):

\[f(x) = (x - 2)(x - 1)(x - 3)\]\[3 \cdot f(x-2) = 3(x - 4)(x-3)(x-5) = (x - 2)(x-1)(x -3 )\]That leaves us with the common of one and all, \(x - 3\)!